# Movable Antenna Enhanced Networked Full-Duplex Integrated Sensing and Communication System

### v2

@@ III-B Optimizing auxiliary variables
$\beta^{\star}_{d,k}={\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{k}}/({{\sum}_{i=1}^{K_{d}}|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{i}|^{2}+\|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{W}}^{r}\|^{2}_{F}+{\sum}_{l=1}^{K_{u}}q_{l}|{h}_{du,k,l}|^{2}+\sigma_{d,k}^{2}}),$ (31)
$\omega^{\star}_{d,k}=\big{(}|1-\beta^{\ast}_{d,k}\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{k}|^{2}+|\beta^{\ast}_{d,k}|^{2}({\sum}_{i\neq k}^{K_{d}}|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{i}|^{2}+\|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{W}}^{r}\|^{2}_{F}+{\sum}_{l=1}^{K_{u}}q_{l}|{h}_{du,k,l}|^{2}+\sigma_{d,k}^{2})\big{)}^{-1},$ (32)
$\beta^{\star}_{u,l}={\sqrt{q_{l}}\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,l}}/({\sum_{i=1}^{K_{u}}q_{i}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,i}|^{2}+\sum_{j=0}^{K_{t}}\sigma_{t,j}^{2}(\sum_{k=1}^{K_{d}}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}|^{2}+\|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}_{r}\|^{2}_{F})+\sigma_{r}^{2}\|\hat{\mathbf{u}}_{l}^{H}\|^{2}_{2}}),$ (33)
$\omega^{\star}_{u,l}=\bigg{(}|1-\beta^{\ast}_{u,l}\sqrt{q_{l}}\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,l}|^{2}+|\beta^{\ast}_{u,l}|^{2}\big{(}{\sum}_{j=0}^{K_{t}}\sigma_{t,j}^{2}({\sum}_{k=1}^{K_{d}}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}|^{2}+\|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}_{r}\|^{2}_{F})+\sigma_{r}^{2}\|\hat{\mathbf{u}}_{l}^{H}\|^{2}_{2}\big{)}\bigg{)}^{-1}$ (34)
@@ III-C Updating The TBS Beamformer
Similarly, the constraint (28b) can also be rewritten as
@@
[…]
@@
It is obviously that the non-convex constraint (43b) makes the problem (P2) intractable. Inspired by the MM framework [20], we respectively linearize the quadratic terms $\hat{\mathbf{w}}_{k}^{H}\mathbf{A}_{9}\hat{\mathbf{w}}_{k}$ and $\bar{\mathbf{w}}^{H}_{r,i}\mathbf{A}_{10,i}\bar{\mathbf{w}}_{r,i}$ as follows
@@
where $\hat{\mathbf{w}}_{k,0}$ and $\bar{\mathbf{w}}^{r}_{i,0}$ are feasible solutions obtained in the last iteration. Therefore, by replace the non-convex terms of constraint (43b) by (41) and (42), the problem (P2) can be rewritten as
[…]
@@
where $\mathbf{a}_{2,k}\triangleq\mathbf{A}_{9}^{H}\hat{\mathbf{w}}_{k,0}$, $\mathbf{a}_{3,i}\triangleq\mathbf{A}_{10,i}^{H}\bar{\mathbf{w}}^{r}_{i,0}$, $\hat{c}_{1,3}\triangleq\!-\!{\sum}_{k=1}^{K_{d}}\!(\hat{\mathbf{w}}_{k,0}^{H}\mathbf{A}_{9}\hat{\mathbf{w}}_{k,0})^{\ast}\!-\!{\sum}_{i=1}^{M_{t}}\!((\bar{\mathbf{w}}^{r}_{i,0})^{H}\mathbf{A}_{10,i}\bar{\mathbf{w}}^{r}_{i,0})^{\ast}+{c}_{1,3}$. The problem (P3) is a typical second order cone program (SOCP) and can be efficiently solved by standard numerical solver, such as CVX [25].
@@ III-E Optimizing The UL User Receiver Filter
With given other variables, the UL user receive filters $\{\hat{\mathbf{u}}_{l}\}$ optimization reduces to solving the following problem
[…]
@@
Since the variables $\{\hat{\mathbf{u}}_{l}\}$ are not intricately coupled in (46a), the problem (P5) can be decomposed into $K_{u}$ independent sub-problems, which each subproblem can be written as
[…]
@@ III-F Optimizing The Sensing Filter
[…]
@@
[…]
The above problem (P8) is typical Rayleigh quotient maximization. By applying the Rayleigh-Rits theorem, the optimal $\hat{\mathbf{u}}_{0}$ can be obtained by aligning with the eigenvector associated with the largest eigenvalue of the matrix $\mathbf{E}_{2}^{-1}\mathbf{E}_{1}$ [27].
Note that although maximizing (P8) does not directly impact the objective value, it will find a more “feasible” variable $\hat{\mathbf{u}}_{0}$, which can provide a larger margin to satisfy the constraint (50b) and hence benefit the optimization of other variables.
@@ III-G Optimizing The TBS MAs’ Position
In this subsection, we discuss updating the position of $m$-th TBS’ $n$-th MA, i.e., $\mathbf{t}^{0}_{m,n}$, when other variables are given. Firstly, the function $\tilde{\mathrm{R}}_{d,k}$, which is defined in (29), can be rewritten as
$-\tilde{\mathrm{R}}_{d,k}=(\mathbf{h}^{0}_{m,k,n})^{H}\mathbf{F}_{5,m,k,n,n}\mathbf{h}^{0}_{m,k,n}$ (53)
$+\textrm{Re}\{(\mathbf{f}^{7}_{m,k,n})^{H}\mathbf{h}^{0}_{m,k,n}\}+c_{4,1,m,k,n},$
where $c_{4,1,m,k,n}$ is a constant and the above newly introduced coefficients are given as
$\mathbf{F}^{1}_{m,n,k,i}\triangleq(\mathbf{w}_{m,i}\mathbf{w}_{n,i}^{H})^{T}\otimes((\mathbf{\Sigma}^{0}_{n,k})^{H}\mathbf{h}^{0,0}_{n,k}(\mathbf{h}^{0,0}_{m,k})^{H}\mathbf{\Sigma}^{0}_{n,k}),$ (54)
$\tilde{\mathbf{h}}^{0}_{m,k}\triangleq\textrm{vec}({\mathbf{H}^{0}_{m,k}}),\mathbf{f}_{m,k,j}^{1}\triangleq{\sum}_{j\neq m}^{M_{t}}(\mathbf{F}^{1}_{m,j,k,i})^{H}\tilde{\mathbf{h}}^{0}_{j,k},$
$\mathbf{F}^{2}_{m,k}\triangleq(\bar{\mathbf{w}}^{r}_{m}(\bar{\mathbf{w}}^{r}_{m})^{H})^{T}\otimes((\mathbf{\Sigma}^{0}_{m,k})^{H}\mathbf{h}^{0,0}_{m,k}(\mathbf{h}^{0,0}_{m,k})^{H}\mathbf{\Sigma}^{0}_{n,k}),$
$\mathbf{f}^{2}_{m,k}\!\!\triangleq\!\!2\omega_{d,k}\beta_{d,k}(\mathbf{w}_{m,k}^{T}\!\!\otimes\!\!(\!(\mathbf{h}^{0,0}_{m,k})^{H}\!\mathbf{\Sigma}^{0}_{m,k})\!)^{H},\mathbf{F}^{4}_{m,k}\!\!\triangleq\!\omega_{d,k}\!|\!\beta_{d,k}\!|^{2}\!\mathbf{F}^{2}_{m,k},$
$\mathbf{F}^{3}_{m,m,k}\!\!\triangleq\!\!\!{\sum}_{i=1}^{K_{d}}\omega_{d,k}|\!\beta_{d,k}\!|^{2}\mathbf{F}^{1}_{m,m,k,i},\mathbf{f}^{4}_{m,k}\!\!\triangleq\!\!\!{\sum}_{i=1}^{K_{d}}\!2\omega_{d,k}|\!\beta_{d,k}\!|^{2}\!\mathbf{f}_{m,k,i}^{1},$
$\mathbf{F}^{5}_{m,k}\triangleq\mathbf{F}^{3}_{m,k}+\mathbf{F}^{4}_{m,k},\mathbf{f}^{6}_{m,k,n}\triangleq 2({\sum}_{i\neq n}^{N_{t}}(\bar{\mathbf{F}}^{5}_{m,k,i,n})^{H}\acute{\mathbf{h}}^{0}_{m,k,i}),$
$\mathbf{f}^{5}_{m,k}\triangleq\mathbf{f}^{4}_{m,k}-\mathbf{f}^{3}_{m,k},\acute{\mathbf{h}}^{0}_{m,k,n}\triangleq\tilde{\mathbf{h}}^{0}_{m,k}[(n-1)L^{t,0}_{m,k}\!+\!1:nL^{t,0}_{m,k}],$
$\bar{\mathbf{F}}^{5}_{m,k,i,j}\!\triangleq\!\mathbf{F}^{5}_{m,k}[(i-1)L^{t,0}_{m,k}\!+\!1\!:\!iL^{t,0}_{m,k},(j-1)L^{t,0}_{m,k}\!+\!1\!:\!jL^{t,0}_{m,k}],$
$\mathbf{f}^{7}_{m,k,n}\triangleq\mathbf{f}^{6}_{m,k,n}+\mathbf{f}^{5}_{m,k}[(n-1)L^{t,0}_{m,k}+1:nL^{t,0}_{m,k}].$
@@
${\sum}_{l=1}^{K_{u}}\tilde{\mathrm{R}}_{u,l}\!=\!\!\!{\sum}_{j=0}^{K_{t}}\!\!-\!\!\textrm{Re}\{(\mathbf{f}^{10}_{m,j,n})^{\ast}\mathbf{g}_{t,m,j}[n]\}\!\!+\!c_{4,2,m,n},$ (55)
where $c_{4,2,m,n}$ is a constant term and the new parameters in (55) being defined as follows
@@
$\mathbf{F}^{6}_{l,j,k}\!\!\triangleq\!\!(\hat{\mathbf{g}}_{r,j}^{H}\hat{\mathbf{u}}_{l}\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\!)^{T}\!\!\!\otimes\!\!(\!\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}\!),\mathbf{f}^{8}_{m,j}\!\triangleq\!\!\!{\sum}_{i\neq m}^{M_{t}}\!\!2(\!\bar{\mathbf{F}}^{9}_{j,m,n}\!)^{H}\!\mathbf{g}_{t,i,j},$
@@
$f^{9}_{m,j,n}\!\!\triangleq\!\!\!{\sum}_{i\neq n}^{N_{t}}\!(\bar{\mathbf{F}}^{9}_{j,m,m}[i,n])^{\ast}\mathbf{g}_{t,m,j}[i],f^{10}_{m,j,n}\!\triangleq\!f^{9}_{m,j,n}\!+\!\mathbf{f}^{8}_{m,j}[n].$

### v3

@@ III-B Optimizing auxiliary variables
$\beta^{\star}_{d,k}={\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{k}}/({{\sum}_{i=1}^{K_{d}}|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{i}|^{2}+\|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{W}}^{r}\|^{2}_{2}+{\sum}_{l=1}^{K_{u}}q_{l}|{h}_{du,k,l}|^{2}+\sigma_{d,k}^{2}}),$ (33)
$\omega^{\star}_{d,k}=\big(|1-\beta^{\ast}_{d,k}\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{k}|^{2}+|\beta^{\ast}_{d,k}|^{2}({\sum}_{i\neq k}^{K_{d}}|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{i}|^{2}+\|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{W}}^{r}\|^{2}_{2}+{\sum}_{l=1}^{K_{u}}q_{l}|{h}_{du,k,l}|^{2}+\sigma_{d,k}^{2})\big)^{-1},$ (34)
$\beta^{\star}_{u,l}={\sqrt{q_{l}}\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,l}}/({{\sum}_{i=1}^{K_{u}}q_{i}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,i}|^{2}+{\sum}_{j=0}^{K_{t}}\sigma_{t,j}^{2}({\sum}_{k=1}^{K_{d}}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}|^{2}+\|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}_{r}\|^{2}_{2})+\sigma_{r}^{2}\|\hat{\mathbf{u}}_{l}^{H}\|^{2}_{2}}),$ (35)
$\omega^{\star}_{u,l}=\bigg(|1-\beta^{\ast}_{u,l}\sqrt{q_{l}}\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,l}|^{2}+|\beta^{\ast}_{u,l}|^{2}\big({\sum}_{j=0}^{K_{t}}\sigma_{t,j}^{2}({\sum}_{k=1}^{K_{d}}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}|^{2}+\|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}_{r}\|^{2}_{2})+\sigma_{r}^{2}\|\hat{\mathbf{u}}_{l}^{H}\|^{2}_{2}\big)\bigg)^{-1}$ (36)
@@ III-C Updating The TBS Beamformer
Similarly, the constraint (30b) can be rewritten as
@@
$\textrm{(P2)}:$ $\mathop{\textrm{min}}\limits_{\{\hat{\mathbf{w}}_{k},\bar{\mathbf{w}}_{r,i}\}}\ {\sum}_{k=1}^{K_{d}}(\hat{\mathbf{w}}^{H}_{k}\mathbf{A}_{5}\hat{\mathbf{w}}_{k}-2\textrm{Re}\{\mathbf{a}^{H}_{1,k}\hat{\mathbf{w}}_{k}\})$ (42a)
@@
It is obvious that the non-convex constraint (45b) makes the problem (P2) intractable. Inspired by the MM framework [35], we respectively linearize the quadratic terms $\hat{\mathbf{w}}_{k}^{H}\mathbf{A}_{9}\hat{\mathbf{w}}_{k}$ and $\bar{\mathbf{w}}^{H}_{r,i}\mathbf{A}_{10,i}\bar{\mathbf{w}}_{r,i}$ as follows
@@
where $\hat{\mathbf{w}}_{k,0}$ and $\bar{\mathbf{w}}^{r}_{i,0}$ are feasible solutions obtained in the last iteration. Therefore, by replacing the non-convex terms of constraint (45b) by (43) and (44), the problem (P2) can be rewritten as
$\textrm{(P3)}:$ $\mathop{\textrm{min}}\limits_{\{\hat{\mathbf{w}}_{k},\bar{\mathbf{w}}_{r,i}\}}\ {\sum}_{k=1}^{K_{d}}(\hat{\mathbf{w}}^{H}_{k}\mathbf{A}_{5}\hat{\mathbf{w}}_{k}-2\textrm{Re}\{\mathbf{a}^{H}_{1,k}\hat{\mathbf{w}}_{k}\})$ (45a)
@@
where $\mathbf{a}_{2,k}\triangleq\mathbf{A}_{9}^{H}\hat{\mathbf{w}}_{k,0}$, $\mathbf{a}_{3,i}\triangleq\mathbf{A}_{10,i}^{H}\bar{\mathbf{w}}^{r}_{i,0}$, $\hat{c}_{1,3}\triangleq\!-\!{\sum}_{k=1}^{K_{d}}\!(\hat{\mathbf{w}}_{k,0}^{H}\mathbf{A}_{9}\hat{\mathbf{w}}_{k,0})^{\ast}\!-\!{\sum}_{i=1}^{M_{t}}\!((\bar{\mathbf{w}}^{r}_{i,0})^{H}\mathbf{A}_{10,i}\bar{\mathbf{w}}^{r}_{i,0})^{\ast}+{c}_{1,3}$. The problem (P3) is a typical second order cone program (SOCP) and can be efficiently solved by standard numerical solvers, such as CVX [42].
@@ III-E Optimizing The UL User Receiver Filter
With the given other variables, the UL user receive filters $\{\hat{\mathbf{u}}_{l}\}$ optimization reduces to solving the following problem
$\textrm{(P5)}:$ $\mathop{\textrm{min}}\limits_{\{\hat{\mathbf{u}}_{l}\}}\ {\sum}_{l=1}^{K_{u}}\{\hat{\mathbf{u}}_{l}^{H}\mathbf{D}_{l}\hat{\mathbf{u}}_{l}-2\textrm{Re}\{\hat{\mathbf{u}}_{l}^{H}\mathbf{d}_{1,l}\}\}+c_{3,1}$ (48a)
@@
Since the variables $\{\hat{\mathbf{u}}_{l}\}$ are not intricately coupled in (48a), the problem (P5) can be decomposed into $K_{u}$ independent subproblems, which each subproblem can be written as
$\textrm{(P6${}_{l}$)}:$ $\mathop{\textrm{min}}\limits_{\hat{\mathbf{u}}_{l}}\ \hat{\mathbf{u}}_{l}^{H}\mathbf{D}_{l}\hat{\mathbf{u}}_{l}-2\textrm{Re}\{\hat{\mathbf{u}}_{l}^{H}\mathbf{d}_{1,l}\}$ (50a)
@@ III-F Optimizing The Sensing Filter
$\textrm{(P7)}:$ $\mathop{\textrm{Find}}\limits_{\hat{\mathbf{u}}_{0}}\ \hat{\mathbf{u}}_{0}$ (52a)
$\textrm{s.t.}\ \frac{\hat{\mathbf{u}}_{0}^{H}\mathbf{E}_{1}\hat{\mathbf{u}}_{0}}{\hat{\mathbf{u}}_{0}^{H}\mathbf{E}_{2}\hat{\mathbf{u}}_{0}}\geq\Gamma,$ (52b)
@@
$\textrm{(P8)}:$ $\mathop{\textrm{max}}\limits_{\hat{\mathbf{u}}_{0}}\ \frac{\hat{\mathbf{u}}_{0}^{H}\mathbf{E}_{1}\hat{\mathbf{u}}_{0}}{\hat{\mathbf{u}}_{0}^{H}\mathbf{E}_{2}\hat{\mathbf{u}}_{0}}$ (54a)
The above problem (P8) is a typical Rayleigh quotient maximization. By applying the Rayleigh-Ritz theorem, the optimal $\hat{\mathbf{u}}_{0}$ can be obtained by aligning with the eigenvector associated with the largest eigenvalue of the matrix $\mathbf{E}_{2}^{-1}\mathbf{E}_{1}$ [44].
@@ III-G Optimizing The TBS MAs’ Position
In this subsection, when other variables are given, we discuss updating the position of $m$-th TBS’ $n$-th MA, i.e., $\mathbf{t}^{0}_{m,n}$  [^5] To enhance readability, we have omitted the subscripts, i.e., $m$ and $n$, from the variables and/or coefficients. Specifically, we rewrite: $\mathbf{t}^{0}_{m,n}$ as $\mathbf{t}^{0}$, $\mathbf{h}^{0}_{m,k,n}$ as $\mathbf{h}^{0}_{k}$, and $\mathbf{g}_{t,m,j}[n]$ as ${g}_{t,j}$.. Firstly, the function $\tilde{\mathrm{R}}_{d,k}$, which is defined in (31), can be rewritten as
$\!-\!\tilde{\mathrm{R}}_{d,k}\!\!=\!\!(\mathbf{h}^{0}_{k})^{H}\mathbf{{F}}^{1}_{k}\mathbf{h}^{0}_{k}\!\!+\!\!\textrm{Re}\{(\mathbf{f}^{1}_{k})^{H}\mathbf{h}^{0}_{k}\}\!\!-\!\!\textrm{Re}\{(\mathbf{f}^{2}_{k})^{H}\mathbf{h}^{0}_{k}\}\!\!+\!\!c_{4,1,k},$ (55)
where $c_{4,1,k}$ is a constant and the above new coefficients are given as follows
$\mathbf{\bar{F}}^{1}_{m,n,k,i}\triangleq(\mathbf{w}_{m,i}\mathbf{w}_{n,i}^{H})^{T}\otimes((\mathbf{\Sigma}^{0}_{n,k})^{H}\mathbf{h}^{0,0}_{n,k}(\mathbf{h}^{0,0}_{m,k})^{H}\mathbf{\Sigma}^{0}_{n,k}),$ (56)
$\tilde{\mathbf{h}}^{0}_{m,k}\triangleq\textrm{vec}({\mathbf{H}^{0}_{m,k}}),\mathbf{f}_{m,k,j}^{15}\triangleq{\sum}_{j\neq m}^{M_{t}}(\mathbf{\bar{F}}^{1}_{m,j,k,i})^{H}\tilde{\mathbf{h}}^{0}_{j,k},$
$\mathbf{F}^{2}_{m,k}\triangleq(\bar{\mathbf{w}}^{r}_{m}(\bar{\mathbf{w}}^{r}_{m})^{H})^{T}\otimes((\mathbf{\Sigma}^{0}_{m,k})^{H}\mathbf{h}^{0,0}_{m,k}(\mathbf{h}^{0,0}_{m,k})^{H}\mathbf{\Sigma}^{0}_{m,k}),$
$\mathbf{f}^{16}_{m,k}\!\!\triangleq\!\!2\omega_{d,k}\beta_{d,k}(\mathbf{w}_{m,k}^{T}\!\!\otimes\!\!(\!(\mathbf{h}^{0,0}_{m,k})^{H}\!\mathbf{\Sigma}^{0}_{m,k})\!)^{H},\mathbf{F}^{4}_{m,k}\!\!\triangleq\!\omega_{d,k}\!|\!\beta_{d,k}\!|^{2}\!\mathbf{F}^{2}_{m,k},$
$\mathbf{F}^{3}_{m,k}\!\!\triangleq\!\!\!{\sum}_{i=1}^{K_{d}}\omega_{d,k}|\!\beta_{d,k}\!|^{2}\mathbf{\bar{F}}^{1}_{m,m,k,i},\mathbf{f}^{17}_{m,k}\!\!\triangleq\!\!\!{\sum}_{i=1}^{K_{d}}\!2\omega_{d,k}|\!\beta_{d,k}\!|^{2}\!\mathbf{f}_{m,k,i}^{15},$
$\mathbf{F}^{5}_{m,k}\triangleq\mathbf{F}^{3}_{m,k}+\mathbf{F}^{4}_{m,k},\mathbf{f}^{2}_{k}\triangleq\mathbf{f}^{16}_{m,k}[(n-1)L^{t,0}_{m,k}+1:nL^{t,0}_{m,k}],$
${\mathbf{F}}^{1}_{m,k,i,j}\!\triangleq\!\mathbf{F}^{5}_{m,k}[(i-1)L^{t,0}_{m,k}\!+\!1\!:\!iL^{t,0}_{m,k},(j-1)L^{t,0}_{m,k}\!+\!1\!:\!jL^{t,0}_{m,k}],$
${\mathbf{F}}^{1}_{k}\!\triangleq\!\mathbf{F}^{5}_{m,k}[(n-1)L^{t,0}_{m,k}\!+\!1\!:\!nL^{t,0}_{m,k},(n-1)L^{t,0}_{m,k}\!+\!1\!:\!nL^{t,0}_{m,k}],$
$\mathbf{f}^{18}_{m,k,n}\triangleq 2({\sum}_{i\neq n}^{N_{t}}({\mathbf{\bar{F}}}^{1}_{m,k,i,n})^{H}\mathbf{h}^{0}_{m,k,i}),$
$\mathbf{f}^{1}_{k}\triangleq\mathbf{f}^{18}_{m,k,n}+\mathbf{f}^{17}_{m,k}[(n-1)L^{t,0}_{m,k}+1:nL^{t,0}_{m,k}].$
@@
${\sum}_{l=1}^{K_{u}}\tilde{\mathrm{R}}_{u,l}\!={\sum}_{j=0}^{K_{t}}-\textrm{Re}\{(f^{3}_{j})^{\ast}{g}_{t,j}\}+c_{4,2},$ (57)
where $c_{4,2}$ is a constant term and the new parameters in (57) are defined as follows
@@
$\mathbf{F}^{6}_{l,j,k}\triangleq(\hat{\mathbf{g}}_{r,j}^{H}\hat{\mathbf{u}}_{l}\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j})^{T}\otimes(\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}),$
@@
$f^{20}_{m,j,n}\triangleq{\sum}_{i\neq n}^{N_{t}}(\bar{\mathbf{F}}^{9}_{j,m,m}[i,n])^{\ast}\mathbf{g}_{t,m,j}[i],$
$\mathbf{f}^{19}_{m,j}\triangleq{\sum}_{i\neq m}^{M_{t}}2(\bar{\mathbf{F}}^{9}_{j,m,n})^{H}\mathbf{g}_{t,i,j},f^{3}_{j}\triangleq f^{20}_{m,j,n}+\mathbf{f}^{19}_{m,j}[n].$